OpenStudy (anonymous):

The non-permissible value of x in the expression: 1/(sqrt(3) - sqrt(x))

OpenStudy (sburchette):

Well there isn't any value that would give us 1/0, so that's not a problem. However, square roots are not defined for numbers less than 0. Because we have sqrt(x), x cannot be less than zero. So any negative number is non-permissible.

OpenStudy (unklerhaukus):

three

OpenStudy (sburchette):

Oh yes, I failed to notice that one, my apologies.

OpenStudy (unklerhaukus):

square roots for number less than zero are complex numbers

OpenStudy (anonymous):

It depends what field you are in. I will suppose you are working with the real numbers... I think you have \[1/(\sqrt{3}-\sqrt{x})\] You cannot divide by 0 (this is undefined. That means you have to solve for when the denominator equals 0. \[\sqrt{3}-\sqrt{x}=0\] \[\sqrt{3}=\sqrt{x}\] Adding sqrt(x) to both sides 3=x (squaring both sides) That means when x=3 we will have 1/0 and so \[x \neq3\] Also if we are working with the real numbers, when we have the square root we cannot have what's inside of it be less than 0 (negative) You have \[\sqrt{x}\] so that means \[x \ge0\] or said another way x cannot be less than or equal to 0. SOLUTION: x cannot equal 3 and x cannot be less than or equal to 0.

OpenStudy (unklerhaukus):

i dont see why complex numbers would be non-permissible

OpenStudy (anonymous):

Thanks guys, I should of specified by the question says "Assuming x is positive"

OpenStudy (anonymous):

IF YOU ARE IN THE FIELD: THE REAL NUMBERS THEN THE COMPLEX NUMBERS WOULD NOT BE PERMISSIBLE. That is why I said "I'll assume/suppose you are working with the real numbers"

OpenStudy (anonymous):

@ Leaper: In that case my answer is applicable except for the part where we have x is greater than or equal to 0, you will now have x is greater than 0. So x cannot be 3 and x cannot be less than 0. Recall that 0 is neither positive or negative. (And i found a mistake above in my answer but after this quick explanation it is void)